OpenStudy (anonymous):

A line passes through the points and . What is the slope of the line that passes through the two points?

OpenStudy (johnweldon1993):

\[slope = \frac{ y_2 - y_1 }{ x_2 - x_1 }\] So you have 2 points (-10 , -4) and (-1 , 2) (x1 , y1) (x2 , y2) plug in those points into that equation so you have \[slope = \frac{ 2 + 4 }{ -1 + 10 }\] What does that simplify to...?

OpenStudy (anonymous):

thank you so much

OpenStudy (johnweldon1993):

No problem!

OpenStudy (anonymous):

A line passes through the points(_10,-4) and (-1,2) What is the y-intercept of the line?

OpenStudy (johnweldon1993):

So for this one....you find the slope (as you just did) and you should have gotten 2/3 This is your slope...now you can put that into the slope intercept formula to solve for your y-intercept.....you can pick either point you have ...lets choose (-1,2) remember this is (x,y) so we have the formula y = mx + b m = slope = 2/3 y = your y-value = 2 x = your x-value = -1 b = your y-intercept so 2 = 2/3(-1) + b Can you solve this for 'b' ...?

OpenStudy (anonymous):

so does b = 1